Where do lines y = x - 1 and y = -5x + 11 intersect?
Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

Hi there,
Solve as simultaneous equations by elimination.
y = x - 1......(1)
y = -5x + 11...(2)
Rearrange
y - x = -1.....(1)
y + 5x = 11....(2)
Multiply (1) by 5
5y - 5x = -5...(1)
y + 5x = 11...(20
Add (1) + (2)
6y.....= 6
y......= 1
Substitute y = 1 into Equatn(1)
y = x - 1
1 = x - 1
x = 2
The lines intersect at (2,1)
Hope this helps :-)

Answer by stanbon(75887)

Where do lines y = x - 1 and y = -5x + 11 intersect?
-------
Substitute and solve for "x"::
x-1 = -5x+11
6x = 12
x = 2
----
Solve for "y"::
y = x - 1
y = 1
-----
Cheers,
Stan H.
------